Mercury cells as “button” batteries are very common and consist of HgO, ZnO, and Zn(OH)2 in a KOH paste. Useful half-cell reactions:

            HgO(s) + H2O(l) + 2e = Hg(l) + 2OH-(aq)              eo = 0.0977 V

            Zn(OH)42-(aq) + 2e = Zn(s) + 4OH-(aq)                 eo = -1.285 V

            Zn(OH)42-(aq) = ZnO(s) + H2O(l) + 2OH-(aq)

A)Such batteries have a + sign marked on them. Which half-reaction is occurring at the electrode so marked?

B)Draw a schematic of a mercury button battery showing the materials making up the anode and cathode.
Answer #1

Mercury batteries use either pure mercury(II) oxide (HgO)—also called mercuric oxide—or a mixture of HgO with manganese dioxide (MnO2) as the cathode. Mercuric oxide is a non-conductor, so some graphite is mixed with it; the graphite also helps prevent collection of mercury into large droplets. The half-reaction at the cathode is:

HgO + H2O + 2e → Hg + 2OH

with a standard potential of +0.0977 V vs. NHE.

The anode is made of zinc (Zn) and separated from the cathode with a layer of paper or other porous material soaked with electrolyte; this is known as a salt bridge. Two half-reactions occur at the anode. The first consists of an electrochemical reaction step:

Zn + 4OH → Zn(OH)4−2 + 2e

followed by the chemical reaction step:

Zn(OH)4−2 → ZnO + 2OH + H2O

yielding an overall anode half-reaction of:

Zn + 2OH → ZnO + H2O + 2e

The overall reaction for the battery is:

Zn + HgO → ZnO + Hg

In other words, during discharge, zinc is oxidized (loses electrons) to become zinc oxide (ZnO) while the mercuric oxide gets reduced (gains electrons) to form elemental mercury. A little extra mercuric oxide is put into the cell to prevent evolution of hydrogen gas at the end of life.

source : Wikipedia

so E0cell = E0reduced - E0oxidesed = 0.0977 - (-1.285) = 1.382 V

At + terminus the annode is situated so recuction takes annode ie...this half reaction

HgO + H2O + 2e → Hg + 2OH

so

cel representation

Zn (s)\left | Zn 2+ (aq) \left | \right | Hg2+ (aq) \left | Hg (s) , E0 cell = 1.382 V